OpenStudy (anonymous):

Would this be no limit or +infinity and -infinity: Find the limit of the function algebraically. limit as x approaches zero of quantity x squared plus two x divided by x to the fourth power.

OpenStudy (anonymous):

\[\lim \frac{ x^2+ 2x}{ x^4 }\]

OpenStudy (anonymous):

x--->0

OpenStudy (anonymous):

same as \(\frac{1}{x^2}+\frac{2}{x^3}\) no limit as you go to zero